OpenStudy (anonymous):

Evaluate 2 x ^ 4 when x equals 3. - ------ x ^ 4

OpenStudy (gazela1997):

2(3)^4 ------ 3^4

OpenStudy (gazela1997):

carry out the equation

OpenStudy (anonymous):

So it would be 2×3=6^4=

OpenStudy (gazela1997):

162 ----- 81

OpenStudy (gazela1997):

and your answer is 2

OpenStudy (anonymous):

ohhhhhhh. I get it

OpenStudy (gazela1997):

np medal mabye

OpenStudy (gazela1997):

?

OpenStudy (anonymous):

Uhhh. How am I supposed to do those

OpenStudy (anonymous):

Do what

OpenStudy (gazela1997):

?

OpenStudy (anonymous):

give medals?

OpenStudy (gazela1997):

you already give me one

OpenStudy (anonymous):

ohhhh . lol